OpenStudy (anonymous):

Find the vertex of 3y=6(x-2)+12

OpenStudy (anonymous):

\[x=\frac{ -b }{ 2a }\]

OpenStudy (anonymous):

first make that a proper equation

OpenStudy (anonymous):

\[3y=6(x-2)+12\] \[3y=6x-12+12\] \[3y=6x\] \[\frac{ 3y=6x }{ 3 }\] \[y=2x\]

OpenStudy (anonymous):

that equation seems wrong there has to be three places. is your equation written correctly?